Ship reports banner with links Home IODP-USIO daily/weekly reports

November >

IODP Expedition 344

Costa Rica Seismogenesis Project,
Program A Stage 2 (CRISP-A2)

Daily Science Report for 31 October 2012

LOCATION: Hole U1380C (8°35.9879′N, 84°4.3918′W, 503 m water depth)

SCIENCE UPDATE: Scientists continued processing discrete samples and working on the Site U1381 reports. The Operations Superintendent gave two ship tours. At Hole U1380C, we continued drilling the 14¾ inch hole for the 10¾ inch casing. After reaching 438 mbsf at 0515 h, we spent the rest of the day preparing the hole for the casing installation. Several sweeps of mud were made to remove fill from the bottom of the hole. While raising the drill bit up to 37 mbsf, excess overpull and torque were experienced from 338 to 396 mbsf, which seems to be the zone with the most instability. When we lowered the bit back to the bottom of the hole, we had to install the top drive at ~280 m to wash and ream from there to total depth. Only 3 m of fill was encountered in the bottom of the hole. While we circulated mud to clean out the fill, we experienced increased torque and the pipe had to be worked free. At the end of the day, we were raising the bit back up with the top drive.


Daily Science Report for 30 October 2012

LOCATION: Hole U1380C (8°35.9879′N, 84°4.3918′W, 503 m water depth)

SCIENCE UPDATE: Scientists turned in the Methods reports for the Proceedings volume and are preparing the Site U1381 reports. A few more samples were taken from Hole U1381C cores. Operations continued at Hole U1380C. The running tool used to deploy the re-entry cone and 16 inch casing was recovered in the early morning hours. To prepare the hole for installing 10¾ inch casing, a 14¾″ drilling assembly was made up and Hole U1380C was reentered at 0550 h. Drilling without recovery started at 0645 h at 48 mbsf and by midnight had reached to 341 mbsf. An additional ~200 styrofoam cups contributed by schools in the San Diego area were lowered to the seafloor on the subsea camera frame.


Daily Science Report for 29 October 2012

LOCATION: Hole U1380C (8°35.9879′N, 84°4.3918′W, 503 m water depth)

SCIENCE UPDATE: Core description was completed for all Hole U1381C cores. The lithology changed at the basal part of the hole from a silty nannofossil calcareous ooze with abundant sponge spicules and foraminifers and rare ash layers to a nannofossil-rich clay with minor glass and feldspar components. In the last meter of the hole, the lithology changed to a consolidated and lithified clay and/or claystone, followed by brecciated oceanic basaltic basement. The nannofossil and radiolaria assemblages suggest that Core U1381C-2H is younger than ~300 ka, Core 6H is younger than 2 Ma, and Core 7H is older than 4 Ma. All core sections have been run through the instrument tracks and measurements of discrete samples continue. All cores have been sampled for postexpedition research. Operations continued at Site U1380. The re-entry cone was lowered through the moon pool at 0855 h and the entire cone and 16-inch casing string assembly were lowered to the seafloor. Hole U1380C was spudded at 1145 h and the re-entry cone landed on the seafloor at 2215 h. About 200 styrofoam cups contributed by schools in the San Diego area were lowered to the seafloor with the subsea camera VIT frame.


Daily Science Report for 28 October 2012

LOCATION: Hole U1380B (8°35.9952′N, 84°4.3908′W, 503 m water depth)

SCIENCE UPDATE: Core description continued for Hole U1381C cores. There is a unit boundary in Core 7H separating hemipelagic green silty clay with nannofossils and terrigenous material above from brown nannofossil-dominated ooze with foraminifers and spicules below. The top of the lower unit is characterized by tephra layers, with 33 identified so far. Measurements were completed on whole-round cores and continued on split cores. Sampling was completed through Core 8H. The jet-in test in Hole U1380B was completed by 1230 h to a depth of 50 mbsf, the ship was offset 20 m to the south over the position of Hole U1380C, and the re-entry cone was moved below the rotary table in the moon pool. The crew spent the rest of the day assembling the 16-inch casing string to be deployed in Hole U1380C, which based on the depth reached by the jet-in test will be 48 m long.


Daily Science Report for 27 October 2012

LOCATION: Hole U1381C (8°25.7027′N, 84°9.4800′W, 2064 m water depth); and
Hole U1380B (8°35.9952′N, 84°4.3908′W, 503 m water depth)

SCIENCE UPDATE: Cores U1381C-5H to 13X advanced from 36.6 to 103.8 mbsf, with a total recovery for Hole U1381C of 109.0 m (105% recovery). Three additional APCT-3 temperature measurements were taken with Cores 5H-7H. Coring at Site U1381 was terminated at 0915 h when the objective of reaching basement was achieved. Measurements on whole-round and split cores continued throughout the day and sampling progressed through the first few cores. The ship was secured for transit and was underway to the next site by 1600 h. The 11-nmi transit was completed in 1.25 h and the ship arrived at Site U1380 at 1715 h. Preparations started for performing a jet-in test in Hole U1380B to determine the length of the first casing string that can be deployed in Hole U1380B.


Daily Science Report for 26 October 2012

LOCATION: In transit from Panama and arrival at Site U1381 (Hole U1381C: 8°25.7027′N, 84°9.4800′W, 2064 m water depth)

SCIENCE UPDATE: The JOIDES Resolution completed the 398-nmi transit from Panama in 36.75 h, at an average speed of 10.8 nmi/h, and arrived at Site U1381 at 0620 h. Scientists completed SampleMaster and DESClogik training, and the Expedition Project Manager gave the final orientations on IODP depth scales and ship reports. A meeting was held with the Siem Offshore crew to discuss the scientific priorities and operations plan of the expedition. The Sample Allocation Committee completed the review of personal sample requests in preparation for Site U1381. The first core from Hole U1381C arrived on deck at 0225 UTC time (2025 local ship time). By midnight we had recovered Cores U1381C-1H to 4H from 0 to 36.6 m below seafloor (mbsf) with a recovery of 38.27 m (105%). An APCT-3 temperature measurement of the formation was taken with Core 3H. Interstitial water and gas analyses started immediately.


Daily Science Report for 25 October 2012

LOCATION: In transit to Site U1381

SCIENCE UPDATE: Scientists attended presentations by the Operations Superintendent on coring technology and the Logging Staff Scientist on CRISP logging measurements. The Curator went over the samples that we anticipate taking on the catwalk and on the sampling table based on shipboard measurements and personal sample requests and some adjustments were made to both. The midnight–noon shift was released at noon to start adjusting to their work shift. The noon–midnight shift was trained on how to use SampleMaster, and the core description and paleontology groups started working with DESClogik. Ship time was retarded by 1 hour to UTC – 6 (U.S. Mountain time zone). Arrival at Site U1381 is expected around 0700 tomorrow morning.


Daily Science Report for 24 October 2012

LOCATION: Rodman Pier 2 North, Panama, and in transit to Site U1381

SCIENCE UPDATE: Scientists attended presentations by Co-Chief Scientist Rob Harris on the CRISP science and operational objectives; the Expedition Project Manager (EPM) on what we need to produce on board and how we are going to accomplish it; the Education Officer and Videographer on education/outreach activities; the Captain, Offshore Installation Manager, and Doctor on ship safety; the Lab Officer and Assistant Lab Officers on life on board and lab safety (including a safety tour); the Marine Computer Specialists on shipboard computer systems; and the EPM and Curator on sampling policy and core flow through the labs (including a tour of the labs and instruments). Lab group meetings were held during the lunch break and as time permitted. The JOIDES Resolution completed fueling and left Rodman Pier at 1800. Following dinner, the core description and paleontology teams were introduced to DESClogik.


Daily Science Report for 23 October 2012

LOCATION: Rodman Pier 2 North, Panama

SCIENCE UPDATE: The Expedition 344 science party welcomed the JOIDES Resolution at Miraflores Locks, Panama. The science party arrived at Rodman Pier around 6 pm and boarded after going through a final immigration check. Scientists spent the evening becoming familiar with the ship.




About the IODP JRSO | Expeditions | Data & core samples | Tools & laboratories | Publications | Education |
Newsroom | Meetings & port calls | Employment | Contact us | Search | Site map | People | Intranet | Home

Accesibility Statement | Plug-in download | Privacy statement | ODP

For comments or questions: webmaster@iodp.tamu.edu

Copyright IODP JRSO

Modified on Wednesday, 12-Dec-2012 13:41:39 CST.